If (9/5)x = (18/5), then x =                                      .
If (9/5)x = (18/5), then x =                                      .
If (9/5)x = (18/5), then x =                                      .

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given equation, (9/5)x = (18/5)
By cross multiplication x = (18/5) × (5/9)
= (18/9)
=2
Then, the value of x = 2

​Thanks